# Alternating Minimization Schemes for Computing Rate-Distortion-Perception Functions with $f$-Divergence Perception Constraints

### v1

@@ Abstract
We study the computation of the rate-distortion-perception function (RDPF) for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the family of $f$-divergences. In this setting, the RDPF forms a convex programming problem for which we characterize the optimal parametric solutions. We employ the developed solutions in an alternating minimization scheme, namely Optimal Alternating Minimization (OAM), for which we provide convergence guarantees. Nevertheless, the OAM scheme does not lead to a direct implementation of a generalized Blahut-Arimoto (BA) type of algorithm due to the presence of implicit equations in the structure of the iteration. To overcome this difficulty, we propose two alternative minimization approaches whose applicability depends on the smoothness of the used perception metric: a Newton-based Alternating Minimization (NAM) scheme, relying on Newton’s root-finding method for the approximation of the optimal iteration solution, and a Relaxed Alternating Minimization (RAM) scheme, based on a relaxation of the OAM iterates. Both schemes are shown, via the derivation of necessary and sufficient conditions, to guarantee convergence to a globally optimal solution. We also provide sufficient conditions on the distortion and the perception constraints which guarantee that the proposed algorithms converge exponentially fast in the number of iteration steps. We corroborate our theoretical results with numerical simulations and draw connections with existing results.
@@ I Introduction
The rate-distortion-perception (RDP) trade-off studies the relevant problem of lossy compression under perceptual constraints on the reconstructed samples, generalizing the classical rate-distortion (RD) analysis [2]. Concurrently proposed by Blau and Michaeli in [3] and Matsumoto in [4, 5], its introduction is motivated by the increasing necessity of a theoretical framework able to incorporate observations raised by a wide body of research spanning from machine learning to multimedia applications, see e.g., [6, 7, 8, 9], remarking the presence of an inherent tension between perceptual quality and fidelity of the compressed samples. In this context, perceptual quality refers to the property of a sample to appear pleasing from the human standpoint. Empirical evaluations of perceptual quality using human scoring [10, 11, 12] show that conventional fidelity measures fail to capture human preference and perception, especially in extreme compression scenarios. Hence, classical rate-distortion theory, focusing solely on the fidelity of the reconstructed samples, may not provide an adequate theoretical framework when considering “human-oriented” data sources, such as images, audio, and video.
The mathematical representation of the RDP trade-off is embodied by the rate-distortion-perception function (RDPF), which complements the classical single-letter rate-distortion function (RDF) with a divergence constraint between the source and reconstruction distributions. The additional constraint acts as a proxy of human perception, measuring the deviation from the real source distribution, also referred to as “natural scene statistic”, following similar principles behind a class of no-reference image quality metrics [13, 14]. It is worth noticing, however, that the selection of specific divergence metrics may be application-dependent and is still an active area of research.
An alternative interpretation of the divergence constraints can be found in its potential as a semantic quality metric, i.e. a quantification of the importance of the reconstructed source from the observer’s perspective [15]. For example, in [16] a comparison between the segmentation capabilities of models trained on traditionally compressed samples against compressed samples with enhanced perceptual quality through Generative Adversarial Network (GAN)-based restoration shows a remarkable improvement in segmentation performances, especially for smaller scene objects.
@@ I-A Related Work
Since its introduction, the RDP trade-off received substantial interest from the information theory community, deriving operational characterizations in a variety of operational scenarios. Focusing on the case where infinite common randomness is available both at the encoder and the decoder, Theis and Wagner [17] provide variable-length codes for both the one-shot and asymptotic regime exploiting the properties of the strong functional representation lemma [18]. In the context of the output-constrained RDF, but also valid for the "perfect realism" RDPF case, Saldi et al. [19] provide coding theorems for the case where only finite common randomness between encoder and decoder is available. In [20], Chen et al. derives coding theorems for the asymptotic case, focusing on the differences between three operational cases defined by the availability of randomness between encoder and decoder, i.e. infinite common randomness, only private randomness, and the deterministic case.
@@
Similar to its classical counterpart, the RDPF does not enjoy any general analytical solution. However, for specific source distributions and particular choices of the distortion and perception measures, there exist closed-form expressions, such as for binary sources subject to Hamming distortion and total variation (TV) distance [3] and Gaussian sources under mean squared-error distortion and various perception measures [21]. The available closed forms, while providing theoretical insights on the RDP trade-off, have limited applicability when considering arbitrary distortion and perception metrics. In the case of the classical RDF, the introduction of the celebrated Blahut-Arimoto (BA) algorithm [22] mitigates the absence of a general closed-form solution, allowing for the exact computation of the RDF for the discrete case. Furthermore, its versatility inspired various approximation methods for the RD in the continuous case [23, 24] and allowed it to be adapted to a variety of source coding problems, with applications from quantum information theory [25] to goal-oriented communication [26]. Also in the RDPF case, numerical solutions have been explored to some extent. Data-driven solutions have been proposed, usually employing a generative adversarial scheme minimizing a linear combination of distortion and perception metrics, see e.g., [3, 27, 28, 29]. Despite the advantage of directly optimizing the image/video codec using only source samples, these methodologies still require particular effort as they are generally highly computational and data-intensive and may suffer from a lack of generalization capabilities. Algorithmic solutions for the RDPF estimation are also present. In the case of discrete alphabets, Chen et al. in [30] cast the RDPF problem as an entropic-regularized Wasserstein barycenter problem and propose a solving method based on the Sinkhorn algorithm applicable to arbitrary distortion measures and perception measure being either a Wasserstein-type distance, the Kullback–Leibler divergence, or the TV distance. Focusing on Gaussian sources, Serra et al. in [21] designs an alternating minimization method for the computation of the Gaussian RDPF for arbitrary fidelity and perception measures, deriving analytical solutions for the so-called “perfect realism” case. Always for the “perfect realism” case, Serra et al. [31] design a solving algorithm for general multivariate continuous sources and distortion metrics, leveraging the information-geometric aspects of the constrained rate-distortion problem. To the best of our knowledge, none of the existing methodologies addresses the generic computation of the RDPF for discrete sources relying on generalizations of the classical BA algorithm.
@@ I-B Contributions
The objective of this work is to propose a generic algorithmic approach for the computation of the RDPF, focusing on the case of discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint that belongs to the class of $f$-divergences.
Our results leverage the fact that the RDPF forms a convex program under mild regularity conditions on the perception constraint (i.e., convexity on the second argument), which are satisfied by the considered class of divergences. This enables us to derive a parametric characterization of the optimal solution of the RDPF (Lemma 1), which is subsequently utilized to construct an alternating minimization procedure, thereon referred to as the Optimal Alternating Minimization (OAM) scheme, for which we also establish convergence guarantees (Theorem 2). However, the resulting structure of the OAM scheme relies on a set of implicit equations in the variables of interest, thus prohibiting the direct implementation of a generic BA algorithm similar to what is already known for the classical rate-distortion theory for $\operatorname{i.i.d.}$ sources and single-letter distortions [22]. Motivated by this technical difficulty, we propose two alternative minimization approaches that solve the implementability issue and whose applicability depends on the smoothness of the considered perception function.
  • In Section III-A, we design a Newton-based Alternating Minimization (NAM) scheme where, noticing that the solution of the OAM iterate is equivalent to a root-finding problem (Lemma 3), we use Newton’s root-finding method [32] for the computation of the optimal iteration step (Theorem 3).
  • In Section III-B, we introduce a Relaxed Alternating Minimization (RAM) scheme, where we leverage a new relaxed formulation of the structure of the iterations in the OAM scheme and, subsequently, we derive necessary and sufficient conditions to ensure convergence to a globally optimal solution (Theorem 4).
In Section IV we design the algorithmic embodiment for the NAM and RAM schemes (see Algorithm 1 and Algorithm 2, respectively) and develop suitable stopping criteria for both algorithms (Theorem 5). Moreover, we also provide sufficient conditions on the structure of the distortion and the perception constraints based on which our algorithms converge exponentially fast in the number of iterations (Theorems 7 and 8). We corroborate our theoretical findings with numerical simulations (Section V), with emphasis on the case of the TV perception metric, for which we develop a smooth approximation (Lemma 8).
@@ I-C Notation
Let $\mathbb{N}$ denote the set of natural numbers, $[a:b]\subset\mathbb{N}$ the integer interval including its extremes, $\mathbb{R}$ the set of real numbers, and $\mathbb{R}_{0}^{+}$ the set of non-negative real numbers. Given a discrete alphabet $\mathcal{X}$, we denote by $(\mathcal{X},\mathbb{B}(\mathcal{X}))$ the Borel measurable space induced by the metric, with $\mathcal{P}(\mathcal{X})$ denoting the set of probability measures defined thereon. We indicate with $\mathcal{Q}(\mathcal{X})$ the set of all transition matrices $Q$ such that $Q\cdot p\in\mathcal{P}(\mathcal{X}^{2})$ for all $p\in\mathcal{P}(\mathcal{X})$. We denote by $\mathbb{E}\left[\cdot\right]$ the expectation operator, and by $\mathbb{E}_{q}\left[\cdot\right]$ we specify the probability distribution $q$ on which the expectation operator is applied. We indicate with square brackets the functional dependency between mathematical objects, e.g. $p[h]$ and $Q[h]$ express the functional dependence of a distribution $p\in\mathcal{P}(\mathcal{X})$ or a transition matrix $Q\in\mathcal{Q}(\mathcal{X})$ on another distribution $h\in\mathcal{P}(\mathcal{X})$. We denote with $C^{n}$ the set of $n^{th}$-times differentiable functions. Given a function $f\in C^{0}$, we denote with $\partial f$ its sub-gradient [33, Definition 8.3], while, if $f\in C^{2}$, we denote by $f^{\prime\prime}(\cdot)$ the second derivative with respect to its argument. We denote by $D(\cdot||\cdot)$ a generic divergence measure whereas $D_{f}(\cdot||\cdot)$ denotes a divergence belonging to the class of $f$-divergences. Given a vector $v\in\mathbb{R}^{d}$, we indicate with $\operatorname{diag}{v}\in\mathbb{R}^{d\times d}$ the matrix with as diagonal the elements of $v$ and zeros otherwise. Given a matrix $V\in\mathbb{R}^{d\times d}$, we denote with $\operatorname{eig}(V)$ its set of eigenvalues.
@@ II Preliminaries
We start this section by providing the formal definition of the RDPF and an overview of its operational meaning, following the works of [3, 17]. Subsequently, given their relevance in this work, we introduce the category of statistical divergences known as $f$-divergences and conclude by providing an overview of the alternating minimization methodology.
@@ II-A Rate-Distortion-Perception Functions
We consider finite alphabet sources and stochastic encoders/decoders pairs having access to a common source of randomness and define the minimum achievable rates when we consider per-letter expected distortion and per-letter perception constraints.
We assume that we are given an $\operatorname{i.i.d.}$ sequence of $n$-length random variables ${X}^{n}\in\mathcal{X}^{n}$ that induce the probability distribution $p_{X}\in\mathcal{P}(\mathcal{X})$. Formally, a stochastic encoder $f^{n}_{E}$ is any function belonging to the set $\mathcal{F}^{n}_{E}=\{f:\mathcal{X}^{n}\times\mathbb{R}\to\mathbb{N}\}$, whereas a stochastic decoder $g^{n}_{D}$ is any function in the set $\mathcal{G}^{n}_{D}=\{g:\mathbb{N}\times\mathbb{R}\to\mathcal{X}^{n}\}$. A stochastic code is an element of $\mathcal{F}^{n}_{E}\times\mathcal{G}^{n}_{D}$. Without loss of generality, the randomness at the encoder and decoder is represented as a single real number, i.e., an infinite number of bits, and is assumed shared by the pair, i.e., common randomness.
We let $d:\mathcal{X}^{2}\to\mathbb{R}_{0}^{+}$ denote a single-letter distortion function and $D:\mathcal{P}(\mathcal{X})^{2}\to\mathbb{R}_{0}^{+}$ denote a divergence measure. Moreover, we define the sets of fidelity criteria $\{\Delta_{i}\}_{i\in[1:n]}$ and $\{\Phi_{i}\}_{i\in[1:n]}$ as follows
@@
where $\Delta_{i}$ is the expected $i^{th}$ symbol distortion and $\Phi_{i}$ is the $i^{th}$ symbol divergence with respect to the reconstructed symbol $Y_{i}$. We are now ready to introduce the definition of achievability and that of the infimum of all achievable rates.
@@
[…]
Next, we give the definition of the information-theoretic characterization of the RDPF [3].
@@
where $D\in[D_{\min},D_{\max}]\subseteq[0,\infty)$, $P\in[P_{\min},P_{\max}]\subseteq[0,\infty)$,
$q_{Y}=\sum_{x\in\mathcal{X}}Q_{Y|X}p_{X}\qquad\text{and}\qquad I(X,Y)=D_{KL}(p_{X}Q_{Y|X}||p_{X}q_{Y})\triangleq I(p_{X},Q_{Y|X})$
@@
We stress the following technical remarks on Definition 2.
@@
(On Definition 2) Following [3], it can be shown that (1) has some useful properties, under mild regularity conditions. In particular, [3, Theorem 1] showed that $R(D,P)$ is (i) monotonically non-increasing function in both $D\in[D_{\min},D_{\max}]\subset[0,\infty)$ and $P\in[P_{\min},P_{\max}]\subset[0,\infty)$; (ii) convex if the divergence $D(\cdot||\cdot)$ is convex in its second argument.
In the sequel, we assume that in (1), the perception constraint is an $f$-divergence, i.e., $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$, which is known to be convex in both arguments [34, Lemma 4.1]. Hence in light of the discussion in Remark 1, $R(D,P)$ forms a convex programming problem.
@@
[…]
Many commonly used divergences functions belong to the class of $f$-divergences. For example,
@@
    $D_{\alpha}(p||q)=\sum_{x\in\mathcal{X}}q(x)f_{\alpha}\left(\frac{p(x)}{q(x)}\right)\qquad$ $f_{\alpha}(x)=\ \begin{cases}\frac{x^{\alpha}-\alpha x-(1-\alpha)}{\alpha(\alpha-1)}\qquad\text{if }\alpha\neq 0,\alpha\neq 1\\ x\log(x)-x+1\qquad\text{if }\alpha=1\\ -\ln(x)+x-1\qquad\text{if }\alpha=0\\ \end{cases}.$
We now provide some general properties of this family of divergences. For any $f$-divergence $D_{f}(\cdot||\cdot)$, the following properties hold:
@@
    $D_{f}(tp_{1}+(1-t)p_{2}||tq_{1}+(1-t)q_{2})\leq tD_{f}(p_{1}||q_{1})+(1-t)D_{f}(p_{2}||q_{2}).$
  • (Invariance) Let $\hat{f}(x)=f(x)+c(x-1)$ for $c\in\mathbb{R}$, then $D_{f}(\cdot||\cdot)=D_{\hat{f}}(\cdot||\cdot)$.
The characterization of the family of $f$-divergences here provided summarizes the properties useful for the scope of this work. For a more in-depth mathematical analysis, we refer the reader to [36].

### v2

@@ Abstract
We study the computation of the rate-distortion-perception function (RDPF) for discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint belonging to the family of $f$-divergences. In this setting, the RDPF forms a convex programming problem for which we characterize optimal parametric solutions. We employ the developed solutions in an alternating minimization scheme, namely Optimal Alternating Minimization (OAM), for which we provide convergence guarantees. Nevertheless, the OAM scheme does not lead to a direct implementation of a generalized Blahut-Arimoto (BA) type of algorithm due to implicit equations in the iteration’s structure. To overcome this difficulty, we propose two alternative minimization approaches whose applicability depends on the smoothness of the used perception metric: a Newton-based Alternating Minimization (NAM) scheme, relying on Newton’s root-finding method for the approximation of the optimal solution of the iteration, and a Relaxed Alternating Minimization (RAM) scheme, based on relaxing the OAM iterates. We show, by deriving necessary and sufficient conditions, that both schemes guarantee convergence to a globally optimal solution. We also provide sufficient conditions on the distortion and perception constraints, which guarantee that the proposed algorithms converge exponentially fast in the number of iteration steps. We corroborate our theoretical results with numerical simulations and establish connections with existing results.
Index Terms:
Lossy Source Coding, Rate-Distortion-Perception Trade-off, Blahut-Arimoto Algorithms, Approximation Methods, $f$-divergence
@@ I Introduction
The rate-distortion-perception (RDP) trade-off studies the relevant problem of lossy compression under perceptual constraints on the reconstructed samples, generalizing the classical rate-distortion (RD) analysis [2]. Concurrently proposed by Blau and Michaeli in [3] and Matsumoto in [4, 5], its introduction is motivated by the increasing necessity of a theoretical framework able to incorporate observations raised by a wide body of research spanning from machine learning to multimedia applications; see e.g., [6, 7, 8, 9], highlighting the presence of an inherent tension between perceptual quality and fidelity of the compressed samples. In this context, perceptual quality refers to the property of a sample to appear pleasing from a human standpoint. Empirical evaluations of perceptual quality using human scoring [10, 11, 12] show that conventional fidelity measures fail to capture human preferences and perceptions, especially in extreme compression scenarios. Hence, classical rate-distortion theory, focusing solely on the fidelity of the reconstructed samples, may not provide an adequate theoretical framework when applied to “human-oriented” data sources, such as images, audio, and video.
The mathematical representation of the RDP trade-off is embodied by the rate-distortion-perception function (RDPF), which complements the classical single-letter rate-distortion function (RDF) with a divergence constraint between the source and reconstruction distributions. The additional constraint acts as a proxy for human perception, measuring the deviation from the real source distribution, also referred to as “natural scene statistic”, following similar principles to those behind a class of no-reference image quality metrics [13, 14]. However, it is worth noting that the selection of specific divergence metrics may be application-dependent and is still an active area of research.
An alternative interpretation of the divergence constraints can be found in its potential as a semantic quality metric, i.e., a quantification of the importance of the reconstructed source to the observer [15]. For example, in [16], a comparison between the segmentation capabilities of models trained on traditionally compressed samples against compressed samples with enhanced perceptual quality through Generative Adversarial Network (GAN)-based restoration shows a remarkable improvement in segmentation performance, especially for smaller scene objects.
@@ I-A Related Work
Since its introduction, the RDP trade-off has received substantial interest from the information theory community, deriving operational characterizations in a variety of operational scenarios. Focusing on the case where infinite common randomness is available both at the encoder and the decoder, Theis and Wagner [17] provide variable-length codes for both the one-shot and asymptotic regime, exploiting the properties of the strong functional representation lemma [18]. In the context of the output-constrained RDF, but also valid for the "perfect realism" RDPF case, Saldi et al. [19] provide coding theorems for the case where only finite common randomness between encoder and decoder is available. In [20], Chen et al. derive coding theorems for the asymptotic case, focusing on the differences between three operational cases defined by the availability of randomness between encoder and decoder, i.e., infinite common randomness, only private randomness, and the deterministic case.
@@
Similar to its classical counterpart, the RDPF does not enjoy any general analytical solution. However, for specific source distributions and particular choices of distortion and perception measures, there exist closed-form expressions, such as for binary sources subject to Hamming distortion and total variation (TV) distance [3] and Gaussian sources under mean squared-error distortion and various perception measures [21, 22, 23]. The available closed forms, while providing theoretical insights into the RDP trade-off, have limited applicability when considering arbitrary distortion and perception metrics. In the case of the classical RDF, the introduction of the celebrated Blahut-Arimoto (BA) algorithm [24] mitigates the absence of a general closed-form solution, allowing for the exact computation of the RDF for the discrete case. Furthermore, its versatility inspired various approximation methods for the RDF in the continuous case [25, 26] and enabled its adaptation to a variety of source coding problems, with applications from quantum information theory [27] to goal-oriented communication [28]. Moreover, in the RDPF case, numerical solutions have been explored to some extent. Data-driven solutions have been proposed, usually employing a generative adversarial scheme minimizing a linear combination of distortion and perception metrics, see, e.g., [3, 22, 29, 30]. Despite the advantage of directly optimizing the image/video codec using only source samples, these methodologies still require considerable effort, as they are generally highly computationally demanding and data-intensive and may suffer from a lack of generalization capabilities. Algorithmic solutions for estimating the RDPF also exist. In the case of discrete alphabets, Chen et al. in [31] cast the RDPF problem as an entropic-regularized Wasserstein barycenter problem and propose a solution method based on the Sinkhorn algorithm applicable to arbitrary distortion measures and with the perception measure being either a Wasserstein-type distance, the Kullback–Leibler divergence, or the TV distance. Focusing on Gaussian sources, Serra et al. in [21] design an alternating minimization method for the computation of the Gaussian RDPF for arbitrary fidelity and perception measures, deriving analytical solutions for the so-called “perfect realism” case. Furthermore, for the “perfect realism” case, Serra et al. [32] design a solution algorithm for general multivariate continuous sources and distortion metrics, leveraging the information-geometric aspects of the constrained rate-distortion problem. To the best of our knowledge, none of the existing methodologies address the generic computation of the RDPF for discrete sources relying on generalizations of the classical BA algorithm.
@@ I-B Contributions
The objective of this work is to propose a generic algorithmic approach for the computation of the RDPF, focusing on the case of discrete memoryless sources subject to a single-letter average distortion constraint and a perception constraint belonging to the class of $f$-divergences.
Our results leverage the fact that the RDPF forms a convex program under mild regularity conditions on the perception constraint (specifically, convexity in the second argument), which are satisfied by the considered class of divergences. This enables us to derive a parametric characterization of the optimal solution of the RDPF (Lemma 1), which is subsequently utilized to construct an alternating minimization procedure, hereafter referred to as the Optimal Alternating Minimization (OAM) scheme, for which we also establish convergence guarantees (Theorem 2). However, the resulting structure of the OAM scheme relies on a set of implicit equations in the variables of interest, thus preventing the direct implementation of a generic BA algorithm, as is already known for the classical rate-distortion theory for $\operatorname{i.i.d.}$ sources and single-letter distortions [24]. Motivated by this technical difficulty, we propose two alternative minimization approaches that address the implementability issue and whose applicability depends on the smoothness of the considered perception function.
  • In Section III-A, we design a Newton-based Alternating Minimization (NAM) scheme observing that the solution of the OAM iterate is equivalent to a root-finding problem (Lemma 3), which allows us to apply Newton’s root-finding method [33] to compute the optimal iteration step (Theorem 3).
  • In Section III-B, we introduce a Relaxed Alternating Minimization (RAM) scheme, where we leverage a new relaxed formulation of the structure of the OAM iterations and subsequently, we derive necessary and sufficient conditions to ensure convergence to a globally optimal solution (Theorem 4).
In Section IV, we design the algorithmic implementations of the NAM and RAM schemes (see Algorithm 1 and Algorithm 2, respectively) and develop suitable stopping criteria for both algorithms (Theorem 5). Moreover, we also provide sufficient conditions on the structure of the distortion and perception constraints under which our algorithms converge exponentially fast in the number of iterations (Theorems 7 and 8). We corroborate our theoretical findings with numerical simulations (Section V), with emphasis on the TV perception metric, for which we develop a smooth approximation (Lemma 8).
@@ I-C Notation
Let $\mathbb{N}$ denote the set of natural numbers, $a:b]\subset\mathbb{N}$ the integer interval including its extremes, $\mathbb{R}$ the set of real numbers, and $\mathbb{R}_{0}^{+}$ the set of non-negative real numbers. Given a discrete alphabet $\mathcal{X}$, we denote by $(\mathcal{X},\mathbb{B}(\mathcal{X}))$ the Borel measurable space induced by the metric, with $\mathcal{P}(\mathcal{X})$ denoting the set of probability measures defined thereon. We indicate with $\mathcal{Q}(\mathcal{X})$ the set of all transition matrices $Q$ such that $Q\cdot p\in\mathcal{P}(\mathcal{X}^{2})$ for all $p\in\mathcal{P}(\mathcal{X})$. We denote by $\mathbb{E}\left[\cdot\right]$ the expectation operator, and by $\mathbb{E}_{q}\left[\cdot\right]$ we specify the probability distribution $q$ on which the expectation operator is applied. We indicate with square brackets the functional dependency between mathematical objects, e.g. $p[h]$ and $Q[h]$ express the functional dependence of a distribution $p\in\mathcal{P}(\mathcal{X})$ or a transition matrix $Q\in\mathcal{Q}(\mathcal{X})$ on another distribution $h\in\mathcal{P}(\mathcal{X})$. We denote with $C^{n}$ the set of $n^{th}$-times differentiable functions. Given a function $f\in C^{0}$, we denote with $\partial f$ its sub-gradient [34, Definition 8.3], while, if $f\in C^{2}$, we denote by $f^{\prime\prime}(\cdot)$ the second derivative with respect to its argument. We denote by $D(\cdot||\cdot)$ a generic divergence measure, whereas $D_{f}(\cdot||\cdot)$ denotes a divergence belonging to the class of $f$-divergences. Given a vector $v\in\mathbb{R}^{d}$, we indicate with $\operatorname{diag}{v}\in\mathbb{R}^{d\times d}$ the matrix with as diagonal the elements of $v$ and zeros otherwise. Given a matrix $V\in\mathbb{R}^{d\times d}$, we denote with $\operatorname{eig}(V)$ its set of eigenvalues.
@@ II Preliminaries
We start this section by providing the formal definition of the RDPF and an overview of its operational meaning, following [3, 17]. Subsequently, given their relevance to this work, we introduce the category of statistical divergences known as $f$-divergences and conclude with an overview of the alternating minimization methodology.
@@ II-A Rate-Distortion-Perception Functions
We consider finite alphabet sources and stochastic encoder/decoder pairs having access to a common source of randomness and define the minimum achievable rates under per-letter expected distortion and per-letter perception constraints.
We assume that we are given an $\operatorname{i.i.d.}$ sequence of $n$-length random variables ${X}^{n}\in\mathcal{X}^{n}$ that induce the probability distribution $p_{X}\in\mathcal{P}(\mathcal{X})$. Formally, a stochastic encoder $f^{n}_{E}$ is any function in the set $\mathcal{F}^{n}_{E}=\{f:\mathcal{X}^{n}\times\mathbb{R}\to\mathbb{N}\}$, whereas a stochastic decoder $g^{n}_{D}$ is any function in the set $\mathcal{G}^{n}_{D}=\{g:\mathbb{N}\times\mathbb{R}\to\mathcal{X}^{n}\}$. A stochastic code is an element of $\mathcal{F}^{n}_{E}\times\mathcal{G}^{n}_{D}$. Without loss of generality, the randomness at the encoder and decoder is modeled as a single real number (i.e., representing an infinite number of bits), and is assumed shared by the pair, i.e., common randomness.
We let $d:\mathcal{X}^{2}\to\mathbb{R}_{0}^{+}$ denote a single-letter distortion function and $D:\mathcal{P}(\mathcal{X})^{2}\to\mathbb{R}_{0}^{+}$ denote a divergence function. Moreover, we define the sets of fidelity criteria $\{\Delta_{i}\}_{i\in[1:n]}$ and $\{\Phi_{i}\}_{i\in[1:n]}$ as follows
@@
where $\Delta_{i}$ is the expected distortion of the $i^{th}$ symbol and $\Phi_{i}$ is the $i^{th}$ symbol divergence with respect to the reconstructed symbol $Y_{i}$. We are now ready to introduce the definition of achievability and that of the infimum of all achievable rates.
@@
$R_{cr}(D,P)\triangleq\inf\{R:R\text{ is $(D,P)$-achievable}\}.$
Next, we state the definition of the information-theoretic characterization of the RDPF [3].
@@
where $D\in[D_{\min},D_{\max}]\subseteq[0,\infty)$, $P\in[P_{\min},P_{\max}]\subseteq[0,\infty)$, $q_{Y}=\sum_{x\in\mathcal{X}}Q_{Y|X}p_{X}$, and
$I(X,Y)=D_{KL}(p_{X}Q_{Y|X}||p_{X}q_{Y})\triangleq I(p_{X},Q_{Y|X})$
@@
In what follows, we highlight certain functional properties of Definition 2.
@@
(On Definition 2 - Functional properties in $(D,P)$) Following [3], it can be shown that (1) has some useful properties, under mild regularity conditions. In particular, [3, Theorem 1] showed that, for $D\in[D_{\min},D_{\max}]\subset[0,\infty)$ and $P\in[P_{\min},P_{\max}]\subset[0,\infty)$, $R(D,P)$ is (i) monotonically non-increasing in both $D$ and $P$; (ii) convex in both $D$ and $P$ if the divergence $D(\cdot||\cdot)$ is convex in its second argument.
Remark 2.
(On Definition 2 - Functional properties in $Q_{Y|X}$) The program defined by (1)-(3) is convex in the transition matrix $Q_{Y|X}$ for a given $p_{X}$ if the divergence $D(\cdot||\cdot)$ is convex in its second argument, since (1) and (2) are respectively convex and affine functions in $Q_{Y|X}$ [35]. Furthermore, the identity kernel $Q_{Y|X}=Id$ always satisfies the constraints given by (2)-(3).
In the sequel, we assume that in (1), the perception constraint is an $f$-divergence, i.e., $D(\cdot||\cdot)=D_{f}(\cdot||\cdot)$, which is known to be convex in both arguments [36, Lemma 4.1].
@@
$\text{(i) }f(0)$ $=\lim_{x\to 0^{+}}f(0),\qquad\text{(ii) }0f\left(\frac{0}{0}\right)=0,$
$\text{(iii) }\forall a\geq 0,~0f\left(\frac{a}{0}\right)=af^{\prime}(\infty).$
Many commonly used divergence functions belong to the class of $f$-divergences. For example,
@@
    $D_{\alpha}(p||q)$ $=\sum_{x\in\mathcal{X}}q(x)f_{\alpha}\left(\frac{p(x)}{q(x)}\right)\qquad$
    $f_{\alpha}(x)$ $=\ \begin{cases}\frac{x^{\alpha}-\alpha x-(1-\alpha)}{\alpha(\alpha-1)}\qquad\text{if }\alpha\neq 0,\alpha\neq 1\\ x\log(x)-x+1\qquad\text{if }\alpha=1\\ -\ln(x)+x-1\qquad\text{if }\alpha=0\\ \end{cases}.$
We now state some general properties of this family of divergences. For any $f$-divergence $D_{f}(\cdot||\cdot)$, the following properties hold:
@@
    $D_{f}(tp_{1}+(1-t)p_{2}||tq_{1}+(1-t)q_{2})$
    $\qquad\qquad\leq tD_{f}(p_{1}||q_{1})+(1-t)D_{f}(p_{2}||q_{2}).$
[…]
The characterization of the family of $f$-divergences provided here summarizes the properties useful for the scope of this work. For a more in-depth mathematical analysis, we refer the reader to [38].